OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \frac{2^{k}k!}{(k+2)!}\] \[\frac{2^{k}}{(k+2)}\] am I on the right track?

OpenStudy (anonymous):

ratio test?

OpenStudy (anonymous):

final answer 4/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{2^{(k+1)}}{2^k} \frac{k+2}{k+3}\]

OpenStudy (anonymous):

i got \[\frac{2(k+1)}{k+3}\] but i could be wrong

OpenStudy (anonymous):

you want to do it slowly?

OpenStudy (anonymous):

I got to that part too then i divided by k

OpenStudy (anonymous):

both numerator and denominator

OpenStudy (anonymous):

? why

OpenStudy (anonymous):

to let k-> \[\infty\]

OpenStudy (anonymous):

i suppose you could but lets think \[\frac{2(k+1)}{k+3}=\frac{2k+6}{k+3}\] numerator is a polynomial of degree 1 as is the denominator since the degrees are the same, the limit as \(k\to \infty\) is the ratio of the leading coefficients, namely 2 i.e. \[\lim_{k\to \infty}\frac{2k+6}{k+3}=\frac{2}{1}=2\] that part you should do in your head from an old pre - calc class

OpenStudy (anonymous):

2(k+1)=(2k +2)

OpenStudy (anonymous):

lol so it is

OpenStudy (anonymous):

of course it make absolutely no difference

OpenStudy (anonymous):

not 2k +6 lolz....so 2/3

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

convergent?

OpenStudy (anonymous):

has nothing at all to do with the constant, just ratio of the leading coefficients

OpenStudy (anonymous):

no divergent, but i have confused you i sense

OpenStudy (anonymous):

so the final answer is 2/3 which is less than 1 so it's convergent

OpenStudy (anonymous):

do you remember a pre calc class of some kind where you looked for horizontal asymptotes?

OpenStudy (anonymous):

if the degree of the numerator is the same as the degree of the denominator, the horizontal asymptote is the ratio of the leading coefficients

OpenStudy (anonymous):

you have \[\lim_{k\to \infty}\frac{2k+2}{k+3}=\frac{2}{1}=2\] so diverges

OpenStudy (anonymous):

forget the constants, they have nothing to do with it

OpenStudy (anonymous):

yes. but here is my point though \[\lim_{k\to \infty}\frac{2k+6}{k+3}=\frac{2}{1}=2\] \[\lim_{k\to \infty}\frac{2k+1}{k+3}=\frac{2}{3}=\frac{2}{3}\]

OpenStudy (anonymous):

\lim_{k\to \infty}\frac{2(k+1)}{k+3}=\frac{2}{3}=\frac{2}{3}

OpenStudy (anonymous):

oh no it has nothing to do with the constant \[\lim_{k\to \infty}\frac{2k+555555}{k+1}=2\]

OpenStudy (anonymous):

\[\lim_{k\to \infty}\frac{2(k+1)}{k+3}=\frac{2}{3}=\frac{2}{3}\]

OpenStudy (anonymous):

the constant is not important in this calculation, it is only the ratio of the leading coefficients

OpenStudy (anonymous):

ok....I must accept that fact then

OpenStudy (anonymous):

this line is incorrect \[\lim_{k\to \infty}\frac{2(k+1)}{k+3}=\frac{2}{3}=\frac{2}{3}\] but you do not have to just accept it, you can make sense out of it as follows

OpenStudy (anonymous):

imagine \(k\) is 1,000,000 then numerator is 2,000,002 and the denominator is 1,000,003

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when you divide, you do not get 2 exactly but it is very very close to 2 right? the constant is unimportant in this calculation because k goes to infinity, i.e. gets very large and the constant is irrelevant

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good! so to finish this problem we see by the ratio test that the limit is 2 and therefore it diverges

OpenStudy (anonymous):

makes sense

OpenStudy (anonymous):

yay